It Took Rakeli and Marisa 30 minutes to skate 5 miles to the ice cream shop. On the trip back home, the route was uphill and it took 75 minutes. Find the average speed for the round trip. Round you answer to the nearest tenth.
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

It Took Rakeli and Marisa 30 minutes to skate 5 miles to the ice cream shop. On the trip back home, the route was uphill and it took 75 minutes. Find the average speed for the round trip. Round you answer to the nearest tenth.
---------------------
The skated 10 miles in 105 minutes, or 1.75 hours.
v = 10miles/1.75 hours = 40/7 mph
v = 5.7 mph
They're not very good skaters.

Answer by checkley77(12844)

D=RT
(5+5)=r(.5+.75)
10=1.25r
10/1.25=r
r=8 mph. was their average speed.